OpenStudy (anonymous):

what is the value of x. x:4 and 45:20 a.8 b.9 c.10

OpenStudy (anonymous):

C is your answer

OpenStudy (anonymous):

OK CAN I ASK YOU ANOTHER

OpenStudy (anonymous):

wait let me recheck

OpenStudy (anonymous):

sorry it was on caps lock

OpenStudy (anonymous):

x/4 = 45/20 cross multiply (20)(x) = (4)(45) 20x = 180 x = 180/20 x = 9

OpenStudy (anonymous):

yup that ^

OpenStudy (anonymous):

21 to x and 7 to 10 a.24 b.30 c.40

OpenStudy (anonymous):

21/x = 7/10 cross multiply (7)(x) = (10)(21) 7x = 210 x = 210/7 x = 30

OpenStudy (anonymous):

ok i have another

OpenStudy (anonymous):

x to 50 and 16 to 25 a.32 b.34 c.41

OpenStudy (anonymous):

x/50 = 16/25 cross multiply (25)(x) = (50)(16) 25x = 800 x = 800/25 x = 32

OpenStudy (anonymous):

ok got another sorry but you are really good at this

OpenStudy (anonymous):

you should try to do one....I will walk you through it

OpenStudy (anonymous):

9/36 and x/4 a.1 b.4 c.9

OpenStudy (anonymous):

look at what kelli does and try to do this one, very easy

OpenStudy (anonymous):

i just need you to give me the answer

OpenStudy (anonymous):

9/36 = x/4 cross multiply (36)(x) = (9)(4) you can finish

OpenStudy (anonymous):

it takes you longer to wait for someone to type it then do it yourself :/

OpenStudy (anonymous):

I just need a million bucks...doesn't mean I am gonna get it

OpenStudy (anonymous):

i have to hurry and get done because i have gymnastics

OpenStudy (anonymous):

answer is : A. 1 Now go do your gymnastics :)

OpenStudy (anonymous):

i have more though

OpenStudy (anonymous):

look at what I have shown you and try to do it yourself. I can no longer do your homework for you. Sorry, but I do not want to get banned

OpenStudy (anonymous):

42/22 and 21/x a.12 b.11 c.10

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

are you mad ?

OpenStudy (anonymous):

i guess not i just dont know this stuff

OpenStudy (anonymous):

do you understand how to cross multiply ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but i will try to look at what you have done

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

 now that is how it is set up